# Digital requirements engineering with an INCOSE-derived SysML meta-model

### v1

@@ 2.2 Structured Requirement
The INCOSE GtWR and Needs and Requirements Manual (NRM) advocates for “structured, natural language” that treats the familiar ‘shall’ statement not as an “atomic entity” but as a “grammatical structure appropriate for communicating needs and requirements” (Wheatcraft et al., 2022). In fact, rule number one (R1) in GtWR is “Structured Statement” which contributes to the quality characteristics: (C3) unambiguous, (C4) complete, (C5) singular, (C7) verifiable, and (C9) conforming (refer to Table 1 and the GtWR Summary Sheet). This INCOSE guidance builds upon a history of success using structured language for textual requirements (Gilb, 2004; Mavin and Wilkinson, 2010; Dick et al., 2017; Carson, 2015, 2021).
Requirement statements with uniform structure are shown to improve quality and have been adopted and recommended by ISO/IEC/IEEE (2018) and INCOSE (2023). Textual “shall” statements with parts in a standard order are easier to write, parse, and verify due to the regular structure that guards against ambiguity and complex grammar. A structured requirement defines a pattern of the requirement statement using placeholders to clearly identify the critical features of the requirement. They have been called template requirements or “statement-level templates” (Wheatcraft and Ryan, 2023), structured requirements (Carson, 2015), requirement structures (Mavin and Wilkinson, 2010), or “boilerplates” (Hull et al., 2011), but the GtWR uses the term ‘requirement pattern’ to avoid confusion with other common uses of the word ‘template’. A requirement pattern is “represented by a series of building blocks (also called pattern slots) including all the elements envisioned to represent a well-formed, singular, and complete requirement” (Wheatcraft and Ryan, 2023). We present the following review of structured requirements using recommended patterns from Carson (2015) and ISO/IEC/IEEE (2018).
Carson-style Requirement Pattern:
@@
“[M]easurable qualitative or quantitative terms specified by characteristics such as an operational scenario, environmental condition, or a cause that is stipulated for a requirement” (Herber et al., 2022).
Carson-style Structured Requirement Example:
@@
The Carson requirement pattern is flexible for use with functional requirements and quality requirements, and the pattern slot names may suit stakeholders who are not as comfortable with SE jargon. Standardizing with this requirement pattern across requirement sets in a project will significantly reduce the ambiguity compared to unstructured requirements. The ISO/IEC/IEEE (2018) standard pattern provides different names for the pattern slots and adds one for [Object] but is otherwise similar to the Carson-style requirement pattern. The standard presents two patterns with an implied “shall” after the [Subject]:
@@
“[M]easurable qualitative or quantitative attributes that are stipulated for a requirement, …and provide attributes that permit a requirement to be formulated and stated in a manner that can be validated and verified” (ISO/IEC/IEEE, 2018).
@@
Many patterns are possible depending on the domain and should be prescribed by the practicing organization (see Appendix C of GtWR for more examples). Here the Subject refers to the part of the system corresponding to the same level as the requirement. The ‘shall’ keyword has been inserted above to clarify the pattern and to emphasize that “requirements are mandatory binding provisions and use ‘shall”’ (ISO/IEC/IEEE, 2018). The Action signifies that the Subject does something — ‘shall not’ is forbidden (R16 in GtWR) — and that the statement is written in the active voice (R2 in GtWR), avoiding superfluous and possibly confusing verbiage such as “be capable of” (R10 in GtWR). This pattern makes clear that every requirement must have a verifiable Constraint. Although every requirement has an associated condition of when it is active, the first pattern may be used in the high-level functional requirements when the Condition is “ubiquitous” (Wheatcraft and Ryan, 2023). The latter pattern may be used as a default value for a redefined ‘Text’ SysML Property. MBSE practice encourages the use of these pattern slots to reference system model elements, not just defined terms in a project glossary, while RE practice as discussed in the GtWR emphasizes the important role of textual requirement statements especially in the presence of system model views used to enhance stakeholder understanding.
2.3 Model-Based Structured Requirement
@@
MBSR is an adaptation of structured requirements to the MBSE paradigm developed by Herber and Eftekhari-Shahroudi (2023) and Herber et al. (2022), building on the pattern concept by making the pattern slots into SysML Properties of a Structured Requirement-stereotyped element (Fig. 4). Thus corresponding model elements and even diagrams may slot-in as model-based supplements to the textual “shall” statement in alignment with the Information-based Needs and Requirements Definition and Management meta-model adopted by the GtWR (Wheatcraft et al., 2019, 2022) (Figs. 5, 6, and 7). This meta-modeling method is well-supported by SysML Profiles and the OMG (2022) specification which describes how SysML requirements can be extended to define requirement types and Property-Based Requirements (Sec. 4.2). MBSR goes “beyond treating a requirement expression as an indivisible entity and allows the terms inside the requirement statement to be referenced” (Dick et al., 2017) as made clear by the unique identifiers in Fig. 3.
⬇ <Model_Based_Structured_Requirements_Profile:Requirement_Expression xmi:id=’_2022x_2_46d01c0_1707158979643_806727_21479_’ base_Class=’_2022x_2_46d01c0_1707158979643_806727_21479’ Id=’L3-EX.1’ Text=’While in the Sample_Collection mode, the Spacecraft shall collect Asteroid_A_Regolith with Regolith_Sample_Mass target between 0.5 kg and 1 kg.’ SR1_Condition=’_2022x_2_46d01c0_1707158979575_932108_21295’ SR2_Subject=’_2022x_2_46d01c0_1707158979604_996809_21373’ SR3_Action=’_2022x_2_46d01c0_1707158979574_907885_21293’ SR4_Object=’_2022x_2_46d01c0_1707158979726_396359_21900’ SR5_Constraint_of_Action=’_2022x_2_46d01c0_1707158979579_344791_21322’ A01_Rationale_Statement_=’Meet the primary mission need’ A08_System_V_V_Primary_Method_=’Test’ A10_System_V_V_Level=’L3-System’ A28_Need_or_Requirement_Verification_Status_=’Complete’ A30_Status_of_the_Need_or_Requirement=’Draft’ A34_Priority_=’High’ A38_Key___Driving=’K+D’ A40_Type_=’Functional’ />
@@
Earlier meta-models of MBSR (Herber et al., 2022) limited the attribute types to corresponding SysML types such as Block for the Subject pattern slot, but this approach was later found to be too restrictive and the standard SysML meta-model decision to use NamedElement was adopted (Herber and Eftekhari-Shahroudi, 2023). The MBSR meta-model works by using Generalization relationships with the existing SysML Requirement, taking advantage of the standard syntax and semantics of SysML Requirements while separating concerns of the Structured Requirement pattern slots, INCOSE GtWR-derived Attributes, and an organization’s conventional requirement attributes. Organization attributes are customizable and might include a secondary V&V method, unique identifier of the associated Work Breakdown Structure unit, or a ‘short text’ which provides an informal requirement statement in layman’s terms (Fig. 5). Other organization attributes may be required for compliance, for syncing with the RMT in use, and for ontological coherence with the ASoT.
@@
The MBSR extensions presented in this paper (Fig. 4) explore the utility of adding the 49 Attributes and 42 Rules described in the INCOSE GtWR with appropriate SysML standard and custom types defined (Figs. 1, 4, and 8). Figure 1 shows a revised meta-model based on the GtWR that clearly relates these DRE terms to each other while respecting their given definitions. The “Requirement Statement” term is replaced with the MBSR “Structured Requirement” and relates Patterns and Pattern Slots; and the “Requirement Expression” is emphasized as the primary element for use so as not to confuse it with the standard SysML Requirement (Sec. 2.1) in the tool interface. A “Requirement Set” contains Requirement Expressions and potentially Requirement Sets, and is a subclass of Requirement Expression rather than a SysML Package to maintain uniform application of the GtWR ontology in the SysML Profile. UML-based multiplicity and roles were used on directed associations to reflect the cross-reference matrices in the GtWR.
The Attributes are grouped according to their purpose and numbered to maintain order
and to aid in searching; some Attributes are marked with an ending asterisk (*) to indicate membership
in the minimum set of Attributes according to the GtWR. Notably, some
Attributes are already defined elsewhere, such as A15 (Unique Identifier) and A16
(Unique Name), and are modeled using
Customization
derived properties that query the
standard SysML properties for completeness of the INCOSE-derived MBSR
profile. A Requirement Set (historically called requirement modules or composite requirements)
is defined to distinguish from an individual Requirement Expression as in the GtWR, and provides
additional querying, filtering, and meta-modeling capabilities for
isolating Sets. Likewise, Needs and Need Sets are defined and inherit
some of the same Attributes, with a different icon to visually
distinguish them from requirements, potentially addressing a major concern with standard SysML voiced in the GtWR.
@@
Value Types such as Enumerations (e.g. Verification Method) and organization-related Stereotypes (e.g. Project Actor, Business Unit) with the Class meta-class were created and selected based on the Attribute definitions and guidance in the NRM. These Value Types and set of minimum viable Attributes should be harmonized with the organization’s conventions, e.g. requirement types, statuses, and risks (Fig. 8).
Attributes help establish Characteristics, of which there are exactly 9
for Needs and Requirements, and 6 for Need Sets and Requirement Sets
(Table 1). When a Requirement Expression or Requirement Set is verified and
validated, a SysML
satisfy
relationship is added from each requirement
to the respective Characteristic, providing model data and metrics of
their well-formedness. Like Attributes, Rules help establish
Characteristics of a well-formed Requirement Expression or Requirement
Set, and during requirements V&V activities, a
satisfy
or new
Violate
relationship is likewise created for metrics and feedback (Fig. 9).
The Rules and Characteristics linked to each requirement may be shown in requirement tables (Fig. 6) and displayed in custom reports.
| ID | Name | Applicability | Derivation | NASA? | ISO? |
| C1 | Necessary | Needs & Requirements | Formal Transformation | ||
| C2 | Appropriate | Needs & Requirements | Formal Transformation | ||
| C3 | Unambiguous | Needs & Requirements | Agreed-to Obligation | ||
| C4 | Complete | Needs & Requirements | Agreed-to Obligation | ||
| C5 | Singular | Needs & Requirements | Formal Transformation | ||
| C6 | Feasible | Needs & Requirements | Agreed-to Obligation | ||
| C7 | Verifiable | Needs & Requirements | Agreed-to Obligation | ||
| C8 | Correct | Needs & Requirements | Formal Transformation | ||
| C9 | Conforming | Needs & Requirements | Formal Transformation | ||
| C15 | Correct | Need Sets & Requirement Sets | Formal Transformation |
@@
A metric suite is provided in the MBSR Profile to demonstrate how validation-based metric definitions and custom scripting can be used to compute completeness metrics on an MBSR set (Herber et al., 2022; The Authors, 2024). The metric suite works by referencing validation rules that check if each MBSR Pattern Slot is filled. The metric table using it shows how many MBSRs are in a given Package at a certain date and time, how many set a value for each Pattern Slot, and what percentage of the MBSRs complete the requirement Pattern. The intended usage is as follows:
@@
The MBSR Completeness metric suite may be used in concert with traceability metric suites to create requirement metrics dashboards. Computing metrics in this way can be more effective than checking every requirement in a matrix, and it provides timestamped data that may be used for burndown charts or compliance audits. See Sec. 5 for a discussion of future work on MBSR metrics.
3 Discussion of Benefits and Challenges
@@
This MBSR SysML Profile (The Authors, 2024) was used to develop over 300 requirements and 50 requirement sets at NASA Jet Propulsion Laboratory (JPL) for the Mars Returned Sample Handling project currently in Pre-Phase A. Drafts of requirements were received from the subsystem engineering leads, imported into No Magic Cameo Systems Modeler 2022x (No Magic, 2023), with given information such as rationale, verification method, verification approach, and additional comments added to the respective GtWR-derived Attributes. Matrix diagrams were created, as in Fig. 9, to aid in the revision of the requirement statements, providing visual feedback as a kind of checklist, and data for metrics used to triage requirements for later revision. All Rules, Attributes, and Characteristics contained the corresponding documentation from GtWR to further assist the usage of the Profile with tooltips. To model Defined Terms as described in GtWR, Cameo Systems Modeler Glossary Tables were filled with Terms and synonyms—often acronyms—active hyperlinks to the definition source, and SysML Allocate relationships to system model elements and diagrams, providing automatic underlining of Defined Terms used consistently throughout requirements statements. In addition, other meta-model elements were added to the Profile to capture the broader range of related information: Goal, Assumption, Project Actor / Role / Organization, Requirement Types relevant to NASA / JPL.

### v2

@@ 2.2 Structured Requirement
The INCOSE GtWR and Needs and Requirements Manual (NRM) advocates for “structured, natural language” that treats the familiar ‘shall’ statement not as an “atomic entity” but as a “grammatical structure appropriate for communicating needs and requirements” (NRM2022). In fact, rule number one (R1) in GtWR is “Structured Statement” which contributes to the quality characteristics: (C3) unambiguous, (C4) complete, (C5) singular, (C7) verifiable, and (C9) conforming (refer to Table 3 and the GtWR Summary Sheet). This INCOSE guidance builds upon a history of success using structured language for textual requirements (gilb2004; mavin2010big; dick2017integrating; carson2015implementing; carson2021).
Requirement statements with uniform structure are shown to improve quality (carson2015implementing; carson2021) and have been adopted and recommended by ISO29148 and INCOSE_SE_HB2023. Textual “shall” statements with parts in a standard order are easier to write, parse, and verify due to the regular structure that guards against ambiguity and complex grammar. A structured requirement defines a pattern of the requirement statement using placeholders to clearly identify the critical features of the requirement. They have been called template requirements or “statement-level templates” (GtWR2023), structured requirements (carson2015implementing), requirement structures (mavin2010big), or “boilerplates” (hull2010requirements), but the GtWR uses the term ‘requirement pattern’ to avoid confusion with other common uses of the word ‘template’. A requirement pattern is “represented by a series of building blocks (also called pattern slots) including all the elements envisioned to represent a well-formed, singular, and complete requirement” (GtWR2023). We present the following review of structured requirements using recommended patterns from carson2015implementing and ISO29148.
Plain-language Requirement Pattern:
@@
“[M]easurable qualitative or quantitative terms specified by characteristics such as an operational scenario, environmental condition, or a cause that is stipulated for a requirement” (herber2022model).
Plain-language Structured Requirement Example:
@@
The plain-language requirement pattern is flexible for use with functional requirements and quality requirements, and the pattern slot names may suit stakeholders who are not as comfortable with SE jargon. Using this requirement pattern consistently across requirement sets will significantly improve readability compared to unstructured requirements. The ISO29148 standard pattern provides different names for the pattern slots and adds one for [Object] but is otherwise similar to the plain-language requirement pattern. The standard presents two patterns with an implied “shall” after the [Subject]:
@@
“[M]easurable qualitative or quantitative attributes that are stipulated for a requirement, …and provide attributes that permit a requirement to be formulated and stated in a manner that can be validated and verified” (ISO29148).
@@
The “measurable outcome” (GtWR2023) that “restrict[s] the design solution or implementation of the systems engineering process” (ISO29148) by applying feasible limits on the [Action] performed by the [Subject] on the [Object] under the stated [Condition].
@@
Many patterns are possible depending on the domain and should be prescribed by the practicing organization (see Appendix C of GtWR for more examples). Here the Subject refers to the part of the system corresponding to the same level as the requirement. The ‘shall’ keyword has been inserted above to clarify the pattern and to emphasize that “requirements are mandatory binding provisions and use ‘shall”’ (ISO29148). The Action signifies that the Subject does something — ‘shall not’ is forbidden (R16 in GtWR) — and that the statement is written in the active voice (R2 in GtWR), avoiding superfluous and possibly confusing verbiage such as “be capable of” (R10 in GtWR). This pattern makes clear that every requirement must have a verifiable Constraint. Although every requirement has an associated condition of when it is active, the first pattern may be used in the high-level functional requirements when the Condition is “ubiquitous” (GtWR2023). The latter pattern may be used as a default value for a redefined ‘Text’ SysML Property. MBSE practice encourages the use of these pattern slots to reference system model elements, not just defined terms in a project glossary, while RE practice as discussed in the GtWR emphasizes the important role of textual requirement statements especially in the presence of system model views used to enhance stakeholder understanding.
2.3 Plain-Language MBSR
@@
MBSR is an adaptation of structured requirements to the MBSE paradigm developed by herber2023 and herber2022model, building on the pattern concept by making the pattern slots into SysML Properties of a Structured Requirement-stereotyped element (Fig. 2). Thus corresponding model elements and even diagrams may slot-in as model-based supplements to the textual “shall” statement in alignment with the Information-based Needs and Requirements Definition and Management meta-model adopted by the GtWR (wheatcraft2019INRDM; NRM2022) (Fig. 3). This meta-modeling method is well-supported by SysML profiles and the SysML1.7 specification which describes how SysML requirements can be extended to define requirement types and Property-Based Requirements (Sec. 5.2). MBSR goes “beyond treating a requirement expression as an indivisible entity and allows the terms inside the requirement statement to be referenced” (dick2017integrating) as commonly done in SysML modeling tools.
@@
Earlier meta-models of MBSR (herber2022model) limited the attribute types to corresponding SysML types such as Block for the Subject pattern slot, but this approach was later found to be too restrictive and the standard SysML meta-model decision to use NamedElement was adopted (herber2023). The MBSR meta-model works by using Generalization relationships with the existing SysML Requirement, taking advantage of the standard syntax and semantics of SysML Requirements while separating concerns of the Structured Requirement pattern slots, and an organization’s conventional requirement attributes. Organization attributes are customizable and might include a secondary system V&V method, unique identifier of the associated Work Breakdown Structure unit, or a ‘short text’ which provides an informal requirement statement in layman’s terms. Other organization attributes may be required for compliance, for syncing with the RMT in use, and for ontological coherence with the ASoT.
3 INCOSE-Derived MBSR for Digital Requirements Engineering
The MBSR extensions presented in this paper (abridged in Fig. 2) explore the utility of adding the 49 Attributes, 42 Rules, and 15 Characteristics described in the INCOSE GtWR with appropriate SysML standard and custom types defined. Extending the prior MBSR profile (Section 2) in this way provided readily-usable SysML classifiers compatible with the INCOSE GtWR ontology, and embedded specific guidance into the SysML modeling environment to aid requirements development and requirements V&V.
3.1 Meta-Model of Rules, Attributes, Characteristics
@@
Figure 4 presents a SysML meta-model derived from the Entity-Relationship Diagram in Figure 4 of GtWR (GtWR2023) that clearly relates these DRE terms to each other while respecting their given definitions. The “Requirement Statement” term is replaced with the MBSR “Structured Requirement” and additionally relates Patterns and Pattern Slots; and the “Requirement Expression” is emphasized as the primary classifier for use so as not to confuse it with the standard SysML Requirement (Sec. 2.1) in the tool’s user interface. A “Requirement Set” contains Requirement Expressions and potentially Requirement Sets, and is a subclass of Requirement Expression rather than a SysML Package to maintain uniform application of the GtWR ontology in the SysML profile. UML-based multiplicity and roles were used on directed associations in Fig. 4 to reflect the cross-reference matrices in the GtWR.
A Requirement Set (historically called requirement modules or composite requirements) is defined to distinguish from an individual Requirement Expression as in the GtWR, and provides additional querying, filtering, and meta-modeling capabilities for isolating Sets. Likewise, Needs and Need Sets are defined and inherit some of the same Attributes, with a different icon to visually distinguish them from requirements, potentially addressing a major concern with standard SysML voiced in the GtWR.
Attributes help establish Characteristics, of which there are exactly 9 for Needs and Requirements, and 6 for Need Sets and Requirement Sets (see Table 3 in Sec. 4.3 for details). The Attributes are grouped according to their purpose and numbered to maintain order and to aid in searching; some Attributes are marked with an ending asterisk (*) to indicate membership in the minimum set of Attributes according to the GtWR. Notably, some Attributes are already defined elsewhere, such as A15 (Unique Identifier) and A16 (Unique Name), and are modeled using «Customization» derived properties that query the standard SysML properties for completeness of the INCOSE-derived MBSR profile (Fig. 11).
@@
Value Types such as Enumerations (e.g. Verification Method) and organization-related Stereotypes (e.g. Project Actor, Business Unit) with the Class meta-class were created and selected based on the Attribute definitions and guidance in the NRM. These Value Types and set of minimum viable Attributes should be harmonized with the organization’s conventions, e.g. requirement types, statuses, and risks (Fig. 5). The Organization Requirement Attributes shown in Fig. 5 are representative, and their modification and use depends on the organization’s context and conventions. This list of attributes should conform to the Organizational Requirement Definition Requirements, as represented in GtWR2023 Figure 7. The presence of such attributes here highlights the flexibility of the MSBR technique while addressing potential concerns that some attributes common in the organization, such as a secondary system V&V method, are missing from the GtWR attributes.
3.2 Requirements V&V using the Profile
We emphasize the distinction between requirements V&V and system V&V, and until the release of the INCOSE Requirements Working Group technical products (GtWR2023; NRM2022), the common systems engineering parlance has been that “requirements verification” refers to system verification against the requirements. Requirements V&V refers to the verification and validation of the requirement expressions themselves, and here we present requirements V&V using the SysML modeling tool and INCOSE-derived MBSR SysML profile.
@@ 3.2 Requirements V&V using the Profile
A metric suite is provided in the MBSR Profile to demonstrate how SysML-validation-based metric definitions and custom scripting can be used to compute completeness metrics on an MBSR set (herber2022model; herber2024repo). The metric suite works by referencing SysML validation rules that check if each MBSR Pattern Slot is filled. The metric table using it shows how many MBSRs are in a given Package at a certain date and time, how many set a value for each Pattern Slot, and what percentage of the MBSRs complete the requirement Pattern. The intended usage is as follows:
@@
The MBSR Completeness metric suite may be used in concert with traceability metric suites to create requirement metrics dashboards. Computing metrics in this way can be more effective than checking every requirement in a matrix, and it provides timestamped data that may be used for burndown charts or compliance audits. See Sec. 6.2 for a discussion of future work on MBSR metrics.
When a Requirement Expression or Requirement Set is verified and validated, a SysML «satisfy» relationship is added from each requirement to the respective Characteristic, providing model data and metrics of their well-formedness. Like Attributes, Rules help establish Characteristics of a well-formed Requirement Expression or Requirement Set, and during requirements V&V activities, a «satisfy» or new «Violate» relationship is likewise created for metrics and feedback (Fig. 6). The Rules and Characteristics linked to each requirement may be shown in requirement tables, metric and generic tables, and rendered in custom reports.
3.3 Application and Case Study
Having presented the INCOSE-derived MBSR profile above (Sec. 3), we now show how the SysML profile and technique are applied using an example in the space domain. The Mars Sample Return project case study is also discussed briefly to explain how the INCOSE-derived MBSR profile was used in a real-world project, starting with unverified draft requirements. Whether the requirements engineer is converting requirement statements to MBSRs or they are developing new requirements in tandem with the system architecture modeling activities, the process is much the same (Fig. 7).
@@ 3.3 Application and Case Study
The INCOSE-derived MBSR SysML profile (herber2024repo) was used to develop over 300 requirements and 50 requirement sets at NASA Jet Propulsion Laboratory (JPL) for the Mars Returned Sample Handling project in Pre-Phase A (Wheaton2025a; Younse2025). Drafts of requirements were received from the subsystem engineering leads in a spreadsheet, and imported into No Magic Cameo Systems Modeler 2022x (Cameo2022x) using the requirements table diagram and the Excel import feature. The given information, such as rationale, verification method, verification approach, and additional comments were first inserted into their respective GtWR-derived Attributes, and continued to be revised during requirements development. Matrix diagrams were created, as in Fig. 6, to aid in the revision of the requirement statements, providing visual feedback as a kind of checklist, and data for metrics used to triage requirements for later revision. All Rules, Attributes, and Characteristics contained the corresponding documentation from GtWR to further assist the usage of the Profile with tooltips. To model Defined Terms as described in GtWR, Cameo Systems Modeler Glossary Tables were filled with Terms and synonyms—often acronyms—active hyperlinks to the definition source, and SysML Allocate relationships to system model elements and diagrams, providing automatic underlining of Defined Terms used consistently throughout requirements statements. In addition, other meta-model elements were added to the Profile to capture the broader range of related information: Goal, Assumption, Project Actor / Role / Organization, Requirement Types relevant to NASA / JPL.
In Fig. 8, we show how a singular INCOSE-derived MBSR may be rendered in a SysML requirements diagram. By showing and hiding compartments and attributes of interest, a custom view is rendered to meet stakeholder concerns. At a glance, a particular requirement expression is shown to be validated according to individual characteristics (Fig. 4), and remaining rules in violation may be shown to draw attention while reviewing the requirement statement. The classical SysML requirement attributes of Id and Text are shown first, followed by categorized attributes numbered accordingly. Lastly, the base MBSR pattern slots are rendered with their associated values’ icons to emphasize the model-based nature of the view. Figure 3.3 shows much of the same information as Fig. 8 but using the standard XML-based data interchange format in support of the tool diversity present in digital engineering environments.
⬇ <Model_Based_Structured_Requirements_Profile:Requirement_Expression xmi:id=’_2022x_2_46d01c0_1707158979643_806727_21479_’ base_Class=’_2022x_2_46d01c0_1707158979643_806727_21479’ Id=’L3-EX.1’ Text=’While in the Sample Collection mode, the Spacecraft shall collect Asteroid A Regolith with Regolith Sample Mass target between 0.5 kg and 1 kg.’ SR1_Condition=’_2022x_2_46d01c0_1707158979575_932108_21295’ SR2_Subject=’_2022x_2_46d01c0_1707158979604_996809_21373’ SR3_Action=’_2022x_2_46d01c0_1707158979574_907885_21293’ SR4_Object=’_2022x_2_46d01c0_1707158979726_396359_21900’ SR5_Constraint_of_Action=’_2022x_2_46d01c0_1707158979579_344791_21322’ A01_Rationale_Statement_=’Meet the primary mission need’ A08_System_V_V_Primary_Method_=’Test’ A10_System_V_V_Level=’L3-System’ A28_Need_or_Requirement_Verification_Status_=’Complete’ A30_Status_of_the_Need_or_Requirement=’Draft’ A34_Priority_=’High’ A38_Key___Driving=’K+D’ A40_Type_=’Functional’ />
While Fig. 8 is shown here more for demonstration purposes, the requirement table is a more common view due to its compactness and ability to sort columns. Figure 10 shows how INCOSE-derived MBSRs may be viewed as a nested tree in a requirements table, with an icon accompanying each model element to again emphasize its model-based nature. Here we also show the numbered characteristics validated, and the numbered rules violated to assist in requirements V&V activities. All structured requirement pattern slots are shown, with some being filled in with “TBD” elements to clearly identify parts of the architecture yet to be modeled and which may be easily searched and enumerated in a separate “TB(X)” table. Attribute “A02-Trace to Parent” is also shown to emphasize traceability of top-level requirements to need expressions in the model. Lastly, a model-based legend is used to clearly identify orphan requirements (red), withdrawn (but not deleted) requirements (gray), and incomplete requirements (yellow). Complete requirements according to their pattern slots are shown in the table with a clear or white background. Such a legend is easily modified using structured expressions and scripts, and, for example, may only show requirements as complete when their minimum attribute set has been filled.
4 Discussion
INCOSE-derived MBSR is a SysML language extension which is leveraged during typical requirements engineering activities, and which is compatible with existing RMT integration. However, compared to an RMT, which supports additional requirement attributes, MBSR validates not only standard data types such as dates and numbers, but also any custom elements created in an architecture model and its accompanying profile. Compared to RMTs, which would require similar customization to include the GtWR rules and characteristics, the MBSR-enabled SysML model makes requirements V&V data directly accessible, supporting a digital engineering approach of data-driven decision-making in a variety of stakeholder views.